### v1

@@ Abstract
Due to the vast array of Android applications, their multifarious functions and intricate behavioral semantics, attackers can adopt various tactics to conceal their genuine attack intentions within legitimate functions. However, numerous feature engineering based methods suffer from a limitation in mining behavioral semantic information, thus impeding the accuracy and efficiency of Android malware detection. Besides, the majority of existing feature engineering based methods are weakly interpretive and fail to furnish researchers with effective and readable detection reports. Inspired by the success of the Large Language Models (LLMs) in natural language understanding, we propose AppPoet, a LLM-assisted multi-view system for Android malware detection. Firstly, AppPoet employs a static method to comprehensively collect application features and formulate various observation views. Subsequently, it steers the LLM to produce function descriptions and behavioral summaries for views via our meticulously devised multi-view prompt engineering technique to realize the deep mining of view semantics. Finally, we collaboratively fuse the multi-view information to efficiently and accurately detect malware through a deep neural network (DNN) classifier and then generate the heuristic diagnostic reports. Experimental results demonstrate that our method achieves a detection accuracy of 97.15% and an F1 score of 97.21%, which is superior to the baseline method Drebin and its variant. Furthermore, the case study evaluates the effectiveness of our generated diagnostic reports.
@@ 1 Introduction
With the advancement of technology, mobile devices have become integral to people’s daily lives. According to "The Mobile Economy 2023" reported by Global System for Mobile communications Association (GSMA) (GSMA (2023)), by the end of 2022, there were 5.4 billion unique mobile subscribers worldwide, with 4.4 billion using mobile internet. By 2030, this number is expected to rise to 6.3 billion for subscribers and 5.5 billion for mobile internet users. Among the range of mobile operating systems, Android has been particularly vulnerable to malware attacks due to its open-source nature. According to statista (statista (2024)), in the third quarter of 2023, over 438,000 instances of mobile malware installation were detected, marking an approximately 19% increase from the second quarter. The proliferation of malware gravely compromises the privacy, property, and personal safety of users, posing significant risks to social stability and national security. Furthermore, as application function continues to expand, malware increasingly seeks to conceal its malicious intent within seemingly legitimate features. Therefore, determining how to effectively detect it remains a persistently pressing issue.
To tackle the issue mentioned above, numerous detection approaches have been proposed. Among them, feature engineering based methods (Arp et al. (2014); Kim et al. (2018); Şahin et al. (2023); Zhu et al. (2023)) have attracted attention for their detection accuracy and generalization ability. These methods extensively extract features from static or dynamic analyses, then select and modify the features for malware identification through machine learning or deep learning algorithms (Akhtar and Feng (2022); Arslan et al. (2019); Kim et al. (2018)). Each target application is transformed to a representation vector to determine its malignancy. Compared with traditional dynamic and static analysis methods (Salehnamadi et al. (2020); Samhi et al. (2022); Enck et al. (2014); Tam et al. (2015)), feature engineering methods have faster processing speed and lower resource consumption in general. Besides, feature engineering provides a customized and easy-to-understand approach. Methods like Graph Neural Network (GNN) that develop rapidly in recent years can automatically learn from graph structures (Hei et al. (2021); Hou et al. (2021); Narayanan et al. (2018)), they are more focused on mining high-dimensional semantic information between features, making it difficult to build intuitive interpretation mechanisms. However, although feature engineering based methods possess a certain degree of interpretability and can identify key factors affecting the discriminatory results to a certain extent, they struggle to produce human-readable and insightful detection reports. In addition, feature engineering based methods are still not sufficiently direct for modeling the explicit semantics between features. Therefore, a mechanism that can better mine behavioral semantic information between features and provide guided detection reports is urgently needed.
Large Language Models (LLMs) have emerged in recent years with impressive performance in natural language understanding tasks. For example, OpenAI’s GPT-3.5 (OpenAI (2024)), trained on massive data resources and with 175 billion parameters, can easily perform a variety of tasks such as text generation (Gao et al. (2023)), language translation (Wang et al. (2023)), program code generation (Jiang et al. (2023)), etc. It has already proved that LLMs have the ability to understand and perform a variety of tasks as knowledgeable experts through its remarkable performance on different domains and topics. Inspired by prompt engineering (Liu et al. (2023a)), this paper leverages the domain expertise accumulated through feature engineering methods to design rigorous and reasonable prompt paradigms for Android malware detection. We employ LLM as the role of an Android security expert to summarize the function of the application features and their potential behaviors in a high-quality text, which allows for further identification and diagnosis. In this way, we can use LLM to further mine the behavioral semantics of features and imporve the interpretability of the detection task.
@@
In this work, we developed AppPoet, an LLM-assisted system for detecting Android malware and generating diagnostic reports. First, AppPoet selects typical features (including permission, API, URL, and uses-feature) accumulated by traditional feature engineering methods (Arp et al. (2014)), and classifies them into different observation views according to the type and information content of the features further. Based on the types of different views, the LLM with rich knowledge is employed to reason and generalize th functions and potential behaviors of the features under each view, thus modeling the behavioral semantics of applications. Subsequently, the pre-trained embedding model is utilized to convert all the textual information into machine-readable representation vectors, which are then fed into the trained DNN classifier (Schmidhuber (2015)) in a multi-view fusion manner to obtain the detection results. By collaborating and complementing the information from different views, AppPoet can improve the performance and generalization of detection. Finally, the LLM is guided to review all known information to generate a readable, valid diagnostic report.
@@
However, a major challenge for AppPoet is to enable LLM to comprehend the features of different views, while outputting the factually correct feature function and inference summary based on the expert knowledge. Specifically, although LLM exhibits strong performance in natural language understanding and logical reasoning, if the task is vaguely defined or overly complex, LLM might process inputs and produce outputs with errors or fabricated content, i.e., LLM’s "hallucination" (Zhang et al. (2023)). Therefore, for the malware detection task, there is a crucial concern about how to involve rich view semantics and numerous types of features, define and decompose the task via effective prompt engineering to ensure LLM outputs align with view semantics and formatting requirements.
@@
To address the above challenge, in this work, we propose the multi-view prompt engineering approach. First, we decompose the task into two phases: function description generation and view summary generation. In the first phase, we provide representative examples to facilitate LLM in generating function descriptions that meet the requirements through in-context learning (Brown et al. (2020); Xie et al. (2021); Work ). Following this, based on the function description lists obtained, we provide LLM with detailed description of the steps, requirements and relevant terminology for generating view summaries by leveraging the concept of chain-of-thought (Wei et al. (2022); Kojima et al. (2022); Wang et al. (2022)), which ensures LLM fully comprehend the task and output the view summary employing the same cognitive process. Subsequently, to accommodate the requirements of multiple views, we meticulously develop function description template and view summary template, which enable LLM to sequentially generate descriptions and summaries for the adapted views by incorporating the feature information from various views. It is worth noting that to further mitigate the negative effects of LLM’s "hallucination," rather than relying solely on LLM to ascertain the maliciousness of a target application based on the known information, our method implements a model cascade approach to train a classifier utilizing a large volume of real samples for the discrimination task. (See Section 4.3 for more details.)
@@
To measure the performance of AppPoet, We conduct a comprehensive evaluation. First, we collect 11,189 real benign apps and 12,128 malicious apps from AndroZoo (Allix et al. (2016)) for training and testing, and our approach achieves higher accuracy, i.e., 97.15% detection accuracy and 97.21% F1 value, compared to the representative feature engineering baseline Drebin and its variant. Second, we evaluate the effectiveness of the multi-view prompt engineering approach used by AppPoet through a comprehensive ablation experiment. In addition, we demonstrate the ability of our fuction memory component to improve efficiency and reduce cost through a comparison experiment. In order to select the most suitable classification model for this work, we also conducted a classification model selection experiment to strengthen AppPoet’s detection capabilities. Finally, we evaluated the instructive value and significance of the diagnostic reports generated by AppPoet through a case. The contributions of this paper are as follows:
@@
  • To the best of our knowledge, our work is an initial exploration of employing LLM for the task of Android malware detection. Based on the powerful inference and summarization capabilities of LLM, we more intuitively generalize the explicit behavioral semantic information among application features to further releasing their detection potential and interpretability.
@@
  • Extensive experiments indicate that our approach outperforms the existing typical baseline Drebin and its variant based on feature engineering. Besides, we validate the effectiveness of the diagnostic reports through a case study.
@@ 2 Related work
In this paper, we are mainly concerned with how to further mine the semantic behavior of the application and improve the interpretability of the detection results based on the features extracted by static methods. Therefore, this section concentrates on detection methods that utilize feature engineering via static methods. Additionally, as our approach is grounded in LLM, we will provide a concise overview of the related concepts and technical applications of LLM.
2.1 Feature engineering based Android malware detection via static methods
Feature engineering based on static methods mainly decompiles APK files through static analysis to obtain AndroidManifest.xml and class.dex, and then further extracts the required features from these files. AndroidManifest.xml contains important configuration information of the Android application, such as package names, permissions (Şahin et al. (2023); Li et al. (2018); Arslan et al. (2019); Arora et al. (2019)), uses-features (Arp et al. (2014)) and component related information (Xu et al. (2016); Cai et al. (2021)), etc. class.dex is the executable file of the application, which can be decompiled into intermediate code (e.g., smali code) or Java source code in the application to recognize and extract the APIs (Onwuzurike et al. (2019); Hou et al. (2016); Qiu et al. (2022)), intents (Xu et al. (2016); Dhalaria and Gandotra (2020)), or control-flow (Yang et al. (2015); Xu et al. (2018)) for use. There are also some works that directly use the binary code of the dex file as a feature (Hsien-De Huang and Kao (2018); Zhu et al. (2023)).
Machine-readable representation vector can be obtained by applying the above features to varying degrees, and then machine learning or deep learning techniques can be used to train classifiers based on large-scale labeled data to accomplish the detection task. Drebin (Arp et al. (2014)) extensively extracted kinds of features within applications, including permissions, APIs, hardware, and network. Then it generated representation vectors based on one-hot coding to detect malware using SVM. Kim et al. (Kim et al. (2018)) used a stacking ensemble learning approach, known as MFDroid, to enhance the detection accuracy by utilizing multiple feature selection algorithms and a logistic regression meta-classifier to handle imbalanced datasets effectively. AppContext (Yang et al. (2015)) used static analysis to extract contextual features of sensitive behaviors in applications, including events that trigger sensitive behaviors and control factors related to sensitive behaviors, and then used machine learning method to distinguish Android applications that contain malicious behaviors. ICCDetector (Xu et al. (2016)) focused on the hidden malicious behaviors in Android’s inter-component and inter-application communication, meticulously analyzes the differences between benign and malicious apps when using Android’s cross component communication related mechanisms, summarizes a set of cross-component communication-related key features and feeds them into a classifier to train a model for malware detection. However, this type of work often only describes the low-level information of the selected features, and are insufficient mining of high-level behavioral semantic information, such as the functions possessed by features and the potential behavioral between features, which can cause false alarms to some extent. Meanwhile, these work are constrained by the fact that black-box models only provide classification results difficult to interpret reliably. Therefore, a mechanism that can further mine high-level semantic information while making effective and readable detection reports based on the information is the direction we are committed to.
@@ 2.2 Large language model
Pre-trained Large Language Models (LLMs) (Chowdhery et al. (2023); Zhang et al. (2022); Ouyang et al. (2022); Touvron et al. (2023); Vaswani et al. (2017)) represented by ChatGPT has opened a new era of natural language understanding, which has been trained on a mega corpus and can support a wide range of natural language processing tasks through prompt engineering Liu et al. (2023a); Chen et al. (2023); Zhou et al. (2022). Compared to NLP with manual design patterns and ML with massive amounts of training data, prompt engineering is extremely lightweight. By simply describing natural language prompts, LLM can be invoked to perform specific tasks without additional training or hard-coding. To further unleash the great potential of LLM, the research community is constantly optimizing prompt engineering methods, such as in-context learning (Brown et al. (2020); Xie et al. (2021); Work ), and chain-of-thought prompting (Wei et al. (2022); Kojima et al. (2022); Wang et al. (2022)).
Due to its powerful text understanding and reasoning capabilities, researchers have employed LLM to solve various tasks within the Android domain Feng and Chen (2024); Liu et al. (2023b, 2024); Huang et al. (2024). To the best of our knowledge, there is still no work that directly uses LLM for Android malware detection, but inspired by LLM’s potential to understand behaviors, the idea of guiding LLM’s inference and summarization through a prompt engineering approach is well suited to be applied to the identification of malicious behaviors.

### v3

@@ Abstract
Due to the vast array of Android applications, their multifarious functions and intricate behavioral semantics, attackers can adopt various tactics to conceal their genuine attack intentions within legitimate functions. However, numerous learning-based methods suffer from a limitation in mining behavioral semantic information, thus impeding the accuracy and efficiency of Android malware detection. Besides, the majority of existing learning-based methods are weakly interpretive and fail to furnish researchers with effective and readable detection reports. Inspired by the success of the Large Language Models (LLMs) in natural language understanding, we propose AppPoet, a LLM-assisted multi-view system for Android malware detection. Firstly, AppPoet employs a static method to comprehensively collect application features and formulate various observation views. Then, using our carefully crafted multi-view prompt templates, it guides the LLM to generate function descriptions and behavioral summaries for each view, enabling deep semantic analysis of the views. Finally, we collaboratively fuse the multi-view information to efficiently and accurately detect malware through a deep neural network (DNN) classifier and then generate the human-readable diagnostic reports. Experimental results demonstrate that our method achieves a detection accuracy of 97.15% and an F1 score of 97.21%, which is superior to the baseline methods. Furthermore, the case study evaluates the effectiveness of our generated diagnostic reports.
@@ 1 Introduction
With the advancement of technology, mobile devices have become integral to people’s daily lives. According to "The Mobile Economy 2023" reported by Global System for Mobile communications Association (GSMA) (GSMA, 2023), by the end of 2022, there were 5.4 billion unique mobile subscribers worldwide, with 4.4 billion using mobile internet. By 2030, this number is expected to rise to 6.3 billion for subscribers and 5.5 billion for mobile internet users. Among the range of mobile operating systems, Android has been particularly vulnerable to malware attacks due to its open-source nature. According to statista (statista, 2024), in the third quarter of 2023, over 438,000 instances of mobile malware installation were detected, marking an approximately 19% increase from the second quarter. The proliferation of malware gravely compromises the privacy, property, and personal safety of users, posing significant risks to social stability and national security. Furthermore, as application function continues to expand, malware increasingly seeks to conceal its malicious intent within seemingly legitimate features. Therefore, determining how to effectively detect it remains a persistently pressing issue.
To tackle the issue mentioned above, numerous detection approaches have been proposed. Among them, learning-based detection methods have attracted attention for their detection accuracy and generalization ability. These methods extract features extensively based on static methods without executing applications. After encoding and transforming the extracted features, a representation vector for each target application is generated, which is then used to train a classifier to distinguish malware from benign applications. Based on how features are utilized, learning-based methods can be classified into three categories: String-based, Image-based, and Graph-based approaches. String-based methods (Arp et al., 2014; Zhu et al., 2023b) primarily arrange the extracted features as sequences of strings, which are then encoded into machine-readable vectors. While these methods are easy to understand and straightforward to implement, they often fail to capture the semantic relationships between features, resulting in decreased detection accuracy. Image-based methods (Sun et al., 2021; Tang et al., 2024) convert APKs into images and apply image recognition techniques for classification. Despite their simplicity and high efficiency, these methods tend to overlook critical semantic information within apps, leading to a reduction in accuracy. Moreover, the use of black-box models complicates result interpretation, limiting the generation of human-readable insights. Graph-based methods (Onwuzurike et al., 2019; Wu et al., 2019; Hou et al., 2021) construct graph structures to capture the semantic relationships between features. Although they can more effectively represent complex application behaviors, constructing large or intricate graph structures introduces challenges related to computational efficiency and resource consumption. Additionally, none of these approaches excel at producing human-readable and insightful reports, making it difficult for security experts to audit and analyze the results effectively.
Large Language Models (LLMs) have recently gained attention for their ability to excel in a wide range of tasks, from natural language understanding to complex reasoning. For example, OpenAI’s GPT-3.5 (OpenAI, 2024), trained on massive data resources and with 175 billion parameters, can easily perform a variety of tasks such as text generation (Gao et al., 2023), language translation (Wang et al., 2023), program code generation (Jiang et al., 2023), etc. These models have proven to be highly versatile, functioning across numerous domains as knowledgeable experts. Leveraging these strengths, we sought to address the limitations of traditional learning-based detection methods by employing LLMs to extract and interpret the semantic relationships within application features. Inspired by LLM’s prompt engineering (Liu et al., 2023a), we designed structured and precise prompt workflows tailored to the characteristics of Android applications. By utilizing LLM to act as Android security analysts, our approach enables the model to analyze feature string sequences, summarize the functions of features, and infer their potential behaviors. Compared to String-based and Image-based methods, our approach leverages LLM to conduct deep semantic analysis of the extracted features. This allows us to capture not only the explicit function meanings but also the implicit relationships between features, thereby improving the accuracy of the analysis. In contrast to Graph-based methods, which can be computationally intensive and difficult to scale, our approach is more efficient and scalable while still maintaining robust interpretability.
@@
In this work, we developed AppPoet, an LLM-assisted system for detecting Android malware and generating diagnostic reports. First, AppPoet selects typical features (including permission, API, URL, and uses-feature) accumulated by traditional String-based methods (Arp et al., 2014), and classifies them into different observation views according to the type and information content of the features further. Based on the different types of views, the LLM, with its rich knowledge, is used to generalize the functions and potential behaviors of the features within each view. Subsequently, the pre-trained embedding model is utilized to convert all the textual information into machine-readable representation vectors, which are then fed into the trained DNN classifier (Schmidhuber, 2015) in a multi-view fusion manner to obtain the detection results. Finally, the LLM is guided to review all known information to generate a readable, valid diagnostic report.
@@
However, a major challenge for AppPoet is to enable LLM to comprehend the features of different views, while outputting the factually correct feature function and inference summary based on the expert knowledge. Specifically, although LLM exhibits strong performance in natural language understanding and logical reasoning, if the task is vaguely defined or overly complex, LLM might process inputs and produce outputs with errors or fabricated content, i.e., LLM’s "hallucination" (Zhang et al., 2023).
@@
To address the above challenge, in this work, we propose the multi-view prompt engineering approach. First, we decompose the task into two phases: function description generation and view summary generation. In the first phase, we provide representative examples to facilitate LLM in generating function descriptions that meet the requirements through in-context learning (Brown et al., 2020; Xie et al., 2021; Work, ). Based on the function description lists, we supply the LLM with detailed steps, requirements, and relevant terminology for generating view summaries using chain-of-thought (Wei et al., 2022; Kojima et al., 2022; Wang et al., 2022) reasoning. This ensures that the LLM fully understands the task and generates summaries with the same cognitive process. To handle multiple views, we carefully designed function and summary templates, allowing the LLM to sequentially generate descriptions and summaries by incorporating feature information from various views. It is worth noting that to further mitigate the negative effects of LLM’s "hallucination," rather than relying solely on LLM to ascertain the maliciousness of a target application based on the known information, our method implements a model cascade approach to train a classifier utilizing a large volume of real samples for the discrimination task.
@@
To assess the performance of AppPoet, we conduct a comprehensive evaluation. First, we collect 11,189 real benign apps and 12,128 malicious apps from AndroZoo (Allix et al., 2016) for training and testing, and our approach achieves higher accuracy, i.e., 97.15% detection accuracy and 97.21% F1 value, compared to the representative learning-based baseline methods (Arp et al., 2014; Onwuzurike et al., 2019; Wu et al., 2019). Second, we evaluate the effectiveness of the multi-view prompt engineering approach used by AppPoet through a comprehensive ablation experiment. In addition, we demonstrate the ability of our fuction memory component to improve efficiency and reduce cost through a comparison experiment. Finally, we evaluated the instructive value and significance of the diagnostic reports generated by AppPoet through a case. The contributions of this paper are as follows:
@@
  • To the best of our knowledge, our work is an initial exploration of employing LLM for the task of Android malware detection. Based on the powerful inference and summarization capabilities of LLM, we intuitively generalize the explicit behavioral semantic information among application features to further releasing their detection potential and interpretability.
@@
  • Experiments indicate that our approach outperforms the existing typical baseline methods. Besides, we validate the effectiveness of the diagnostic reports through a case study.
@@ 2 Related work
2.1 Learning-based Android malware detection
Learning-based Android malware detection methods accomplish detection task mainly through machine learning and deep learning techniques. This type of methods usually starts by decompiling the APK. Next, a variety of different features are extracted and selected, and are combined and processed in different ways to obtain the applied representation vector. Finally, these representation vectors are used to train a classifier to detect malware. We categorize learning-based methods into String-based, Image-based, and Graph-based in terms of the different uses of features and introduce each of them separately.
2.1.1 String-based detection methods
String-based methods typically organize features into a sequence of strings, which are then encoded into machine-readable vectors for training classifiers. Different researchers have selected a wide variety of features to construct sequences from different perspectives. Some works (Li et al., 2018; Arslan et al., 2019; Şahin et al., 2023) get permissions declared by apps from the AndroidManifest.xml files and compose sequences with different feature selection methods. APIs are also one of the key features of interest to the research community. AppContext (Yang et al., 2015) leverage static analysis to extract contextual features of sensitive APIs in apps, including events that trigger sensitive APIs and control factors related to sensitive APIs. Yumlembam et al. (2023) captures the difference in usage between benign and malware APIs by introducing the BM25 (Best Matching 25) scoring feature, which calculates the BM25 score for each API. In addition to focusing on one key feature to construct a sequence, more methods (Arp et al., 2014; Kim et al., 2018; Dhalaria and Gandotra, 2020; Qiu et al., 2022; Cai et al., 2021; Zhu et al., 2023b) use a wide range of features to construct representation sequences. The most representative is Drebin (Arp et al., 2014), which extensively extracted kinds of features within apps, including permissions, APIs, hardware, and network. String-based methods often employ feature engineering to select or refine features, but they lack sufficient depth in semantic analysis. This limitation makes it difficult to recognize contextual associations and interactions among features to uncover potential malicious behavior patterns, which adversely affects detection accuracy.
2.1.2 Image-based detection methods
Image-based methods consider how bytecode can be converted into an image and then detected using image recognition algorithms. These methods (Hsien-De Huang and Kao, 2018; Xiao and Yang, 2019) typically map bytecode to each channel of RGB, thus converting .dex file to an RGB image, and then use these RGB-encoded representation vectors to train CNN classifiers. For better detection, recent work has used more complex and advanced classification algorithms instead of CNNs. Zhu et al. (2023a) selects vital parts of .dex file to be described as an RGB image and then uses the proposed novel CNN variant classifier for detection. Tang et al. (2024) proposes a method based on novel hybrid bytecode image and deep neural network combined with attention mechanism. Sun et al. (2021) combine .dex file, .so file, and .xml file by mapping them to different RGB channels to create a image for detection. It should be noted that such methods tend to use an end-to-end architecture. They can be rapidly used for detection after obtaining features with only a relatively shallow-level of processing. It means that these methods have a huge advantage in terms of detection efficiency, but they ignore critical semantic information in apps thus causing a loss of accuracy. In addition, Image-based methods are often regarded as black-box models that are difficult to interpret.
2.1.3 Graph-based detection methods
Different from processing features as a sequence of strings or converting them to image, Graph-based methods use the extracted features to construct graph structures that contain various semantic information. MaMaDroid (Onwuzurike et al., 2019) statically extracts API calls from APKs and abstracts them into the form of family calls or package calls, and then models the feature vectors for training the classification model through Markov chains. Malscan (Wu et al., 2019) constructs function calls graphs from smali file of APKs and selects the sensitive API calls from it based on PScout. CDGDroid (Xu et al., 2018) uses control-flow graphs, data-flow graphs and their possible combinations as features to characterize APKs. AMCDroid (Liu et al., 2023c) models application behavior as a homogeneous graph based on call graphs and code statements. Chen et al. (2024) proposed a new type of call graph called the class-set call graph (CSCG), which takes Java class sets as nodes and call relationships between class sets as edges. These methods then use different graph representation learning methods to transform these graphs into representation vectors to train classification models for detection. In addition, there also some works (Hei et al., 2021; Hou et al., 2021; Ye et al., 2019) extract a wide range of features from APKs to construct heterogeneous information networks for malware detection. There is no doubt that Graph-based methods are far more capable of mining semantic information for apps than String-based and Graph-based methods. However, they still have the following limitations. On one hand, constructing a graph that is sufficient to adequately represent semantic information consumes lots of resources, especially for methods based on static analysis to obtain information-flow graphs. On the other hand, while graph-based methods implicitly mine the semantics of application behaviors through graph representations, they often struggle to provide intuitive insights for human experts to conduct fine-grained auditing and analysis.
@@ 2.2 Large language model
Pre-trained Large Language Models (LLMs) (Chowdhery et al., 2023; Zhang et al., 2022; Ouyang et al., 2022; Touvron et al., 2023; Vaswani et al., 2017) represented by ChatGPT has opened a new era of natural language understanding, which has been trained on a mega corpus and can support a wide range of natural language processing tasks through prompt engineering (Liu et al., 2023a; Chen et al., 2023; Zhou et al., 2022). Compared to NLP with manual design patterns and ML with massive amounts of training data, prompt engineering is extremely lightweight. By simply describing natural language prompts, LLM can be invoked to perform specific tasks without additional training or hard-coding. To further unleash the great potential of LLM, the research community is constantly optimizing prompt engineering methods, such as in-context learning (Brown et al., 2020; Xie et al., 2021; Work, ), and chain-of-thought prompting (Wei et al., 2022; Kojima et al., 2022; Wang et al., 2022).
Due to its powerful text understanding and reasoning capabilities, researchers have employed LLM to solve various tasks within the Android domain (Feng and Chen, 2024; Liu et al., 2023b, 2024; Huang et al., 2024). To the best of our knowledge, there is still no work that directly uses LLM for Android malware detection, but inspired by LLM’s potential to understand behaviors, the idea of guiding LLM’s inference and summarization through a prompt engineering approach is well suited to be applied to the identification of malicious behaviors.